What would be the resulting pH if 5ml of 10M HCl were added to 1 litre of pure water?
1
Expert's answer
2019-09-09T07:07:33-0400

Solution.

Find the amount of HCl substance in 5 ml with a concentration of 10 M.

"C(HCl) = \frac{n}{V}"

"n = C(HCl) \times V"

n = 0.05 mole

Find the total volume of the solution after mixing 1 liter of water and 5 ml of hydrochloric acid. This is the sum of these volumes.

"V(total) = V(H2O) + V(HCl)"

V(total) = 1.005 liters

Find the concentration of hydrochloric acid in the new solution obtained.

"C'(HCl) = \frac{n}{V(total)}"

C'(HCl) = 0.05 M

Find the pH of the resulting solution. Since hydrochloric acid is a strong acid, we use this formula:

"pH = -lg(C'(HCl))"

pH = 1.30

Answer:

pH = 1.30